OpenStudy (anonymous):

The average (arithmetic mean) of a,b,and c, is equal to the median of a, b,and c. If 0>a>b>c, which of the following must be equal to b?

OpenStudy (anonymous):

A. a+c/2 B. a+c/3 C. c-a/2 D. C-A/3 E. Sqrt AC

OpenStudy (anonymous):

Given 0>a>b>c. b must be the median. The average = (a+b+c)/3. So b= (a+b+c)/3 and solve for b....hint: distribute out the fraction

OpenStudy (anonymous):

So would it be B?

OpenStudy (anonymous):

b = a/3 + b/3 + c/3 (subtract b/3 from both sides) 2/3b = a/3 + c/3 (multiply 3/2 to both sides) b = a/2 + c/2 or b = (a+c) /2

OpenStudy (anonymous):

oh i forgot to factor out the /3 along with b.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and this answer makes sense for b to be the average....the average of a and c (a+c/2) must equal b as well

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

no problem